OpenStudy (anonymous):

2tan((pi/2)-x)sin^(2)x=(sqrt(3)/2) PLEASE Help! :)

OpenStudy (anonymous):

abyss, do you understand this? I am completely lost!

OpenStudy (anonymous):

sin [pi/2 - x] = cos x cos [pi/2 - x] = sin x ---> tan [Pi/2 - x ] = cos x/sin x (2*cos x*sin^2 x)/sin x = 2sin x*cos x = sin 2x = sqr3/2 = sin 2Pi/3 2x = 2Pi/3 --> x = Pi/3 (0 , x < 2Pi)

OpenStudy (anonymous):

Thank you! I forgot about that identity. :)